Angle:

Wilma and Greg were trying to solve the quadratic equation: x^2 + bx + c = 0 Wilma wrote down the wrong value of b (but her value of c was correct), and found the roots to be 1 and 6. Greg wrote down the wrong value of c (but his value of b was correct) and found the roots to be -1 and -4. What are the actual roots of x^2 + bx + c = 0?

Angle:

roots of 1 and 6 mean: (x - 1) (x - 6) = 0 this expands to: x^2 - 7x + 6 = 0 then roots of -1 and -4 mean: (x + 1) (x + 4) = 0 this expands to x^2 + 5x + 4 = 0

Angle:

so we take Wilma's c value and Greg's b value b = 5 c = 6

Angle:

x^2 + 5x + 6 = 0 6 = 2 x 3 2+3 = 5 x^2 + 2x + 3x + 6 = 0 [factor by grouping] x(x+2) + 3(x+2) = 0 (x+2)(x+3) = 0

Angle:

so the roots are -2 and -3